# Gender-Neutral Large Language Models for Medical Applications: Reducing Bias in PubMed Abstracts

### v1

@@ Methods
| Bert-Base (Comparison Baseline for All Cases) | 40% |
| PubMedBert | 20% |
| 1965Bert | 4% |
| MOBERT | 70% |
The MOBERT results were further analyzed to determine a relationship between the frequency of the occupational term in the training data and the accuracy of replacement, as shown in Table 9.
| Occupational Term | Frequency | Percentage |
| Physician | 298 | 100% |
| Surgeon | 135 | 100% |
| Doctor | 89 | 70% |
| Practitioner | 68 | 60% |
| Nurse | 64 | 30% |
Discussion
The application of our gender-neutralization pipeline to the 1965-1980 PubMed abstracts has demonstrated its potential to significantly reduce occupational gender bias in large language models. By introducing gender-neutral pronouns reconciled with occupational terms in 1,400 abstracts, we successfully trained a model, MOBERT, that demonstrated a 70% success rate in predicting inclusive pronouns in a masked language modeling task. This result far exceeds the 4% success rate of 1965Bert, a model trained on unmodified texts from the same period, and highlights the importance of correcting biased data at the training stage. MOBERT’s performance also surpassed that of both the base model, Bert-Base, which exhibited a 40% success rate, and Full PubMedBert, a model trained on the complete PubMed dataset without gender-neutralization, which achieved only a 20% inclusive successive rate. These comparisons underscore the critical role of targeted intervention in mitigating bias in language models.
Our results confirm the hypothesis that training models on biased, gendered data leads to a decline in their ability to generate gender-neutral pronoun predictions for occupations as subjects. The 1965-1980 dataset, which had a notable prevalence of gendered language, resulted in 1965Bert showing a marked decrease in its ability to predict inclusive pronouns. Conversely, MOBERT’s improved performance illustrates that interventions like ours can substantially counteract this bias. This suggests that gender-neutralization efforts in model training can directly enhance a model’s ability to use inclusive language, particularly in domains such as healthcare, where the presence of occupational gender bias could potentially skew diagnoses and treatment recommendations.
Despite the overall success of MOBERT, our analysis did reveal some limitations. As seen in Table 8, the frequency of occupational terms in the dataset correlated strongly with the accuracy of pronoun replacement. For instance, terms like “physician” and “surgeon,” which appeared more frequently in the training data, saw a 100% accuracy in neutral pronoun predictions, while terms like “nurse” had a much lower replacement rate of 30%. This indicates that the more exposure the model had to gendered language associated with a particular occupation, the more effectively it could neutralize pronouns related to that term. This suggests that expanding the dataset to include a broader range of occupations and more balanced representation of male- and female-dominated roles could further improve the model’s performance.
Another potential limitation arises from the process of language alteration itself. Although we carefully designed our pipeline to neutralize pronouns only in contexts where the occupational term was the antecedent, there remains a risk that some instances of gendered language with medically significant context may have been inadvertently modified. While we found no evidence of such errors in our testing, further refinements to the pipeline could include more sophisticated contextual analysis to ensure the protection of patient-specific or trial-related information.
In future work, we plan to expand our study to process the entire PubMed dataset, which will allow us to analyze a wider range of occupational terms and scenarios. This would provide a more comprehensive understanding of how LLMs can be trained to avoid occupational gender bias across a diverse range of medical contexts. Additionally, applying this pipeline to more recent texts would enable us to track changes in gender bias over time and assess whether such biases persist in more contemporary medical literature.
Conclusion
@@
This work demonstrates the effectiveness of a gender-neutralization pipeline in reducing occupational gender bias in large language models trained on medical literature. By processing 379,000 PubMed abstracts from 1965-1980 and targeting gender-specific pronouns linked to professions, we improved MOBERT’s success rate to 70% in predicting gender-neutral pronouns, compared to 4% for 1965Bert. This improvement highlights the importance of addressing bias during training. While promising, the study also reveals opportunities for improvement, such as expanding the dataset and refining the pipeline to ensure complete accuracy. These findings underscore the potential for creating more equitable and unbiased models in medical and other sensitive domains.
@@ Acknowledgements
This work was supported by the Cancer Prevention and Research Institute of Texas (CPRIT RP210045) as part of the Biomedical Informatics, Genomics and Translational Cancer Research Training Program.

### v2

@@ 4.1 Principal Results
| BERT-Base (Comparison Baseline for All Cases) | 40 |
| PubMedBERT | 20 |
| 1965BERT | 4 |
| MOBERT | 70 |
4.2 Comparison with Prior Work
Research on the recruitment and retention of women in male-dominated occupations highlights how deeply embedded gendered language can reinforce exclusionary workplace cultures Germain et al. (2012). Prior studies such as De-Arteaga et al. (2019) have shown that occupational gender bias in machine learning software can directly affect hiring and professional representation. Research has shown that the assumption of male dominance in professional fields can discourage the participation of women in those fields, leading to self-reinforcing cycles of underrepresentation Wu (2022). For instance, studies have demonstrated that increasing gender diversity in male-dominated academic settings leads to improved career outcomes for female students, suggesting that removing implicit assumptions – such as assuming doctors are male – could encourage more diverse participation in medicine, technology, engineering, and math Germain et al. (2012). If gendered language in professional texts perpetuates the underrepresentation of women in these fields, systematically neutralizing such biases could contribute to breaking this cycle.
4.3 Future Improvements
Future studies could involve integrating MOBERT into clinical NLP applications – such as diagnostic models and medical literature retrieval systems – to assess whether gender neutralization leads to improved healthcare equity. Sun et al. (2019) highlight that debiasing techniques should be evaluated not only through linguistic performance but also through real-world impact within medicine. Conducting user studies with medical professionals would be useful in assessing how gender-neutral models influence literature search relevance and clinical decision-making.
| Occupational Term | Frequency | Percentage (%) |
| Physician | 298 | 100 |
| Surgeon | 135 | 100 |
| Doctor | 89 | 70 |
| Practitioner | 68 | 60 |
| Nurse | 64 | 30 |
5 Conclusions
@@
This work demonstrates the effectiveness of a gender neutralization pipeline in reducing occupational gender bias in large language models trained on medical literature. By processing 379,000 PubMed abstracts from 1965-1980 and targeting gender-specific pronouns linked to professions, we improved MOBERT’s success rate to 70% in predicting gender-neutral pronouns, compared to 4% for 1965BERT. This improvement highlights the importance of addressing bias during training. While promising, the study also reveals opportunities for improvement, such as expanding the dataset and integrating the pipeline into further applications. These findings underscore the potential for creating more equitable and unbiased models in medical and other sensitive domains.
Limitations
Despite the overall success of MOBERT, our analysis did reveal some limitations. As seen in Table 8, the frequency of occupational terms in the 379,000 modified abstracts used for pre-training correlated strongly with the accuracy of pronoun replacement. For instance, terms like “physician” and “surgeon,” which appeared more frequently in the training data, saw a 100% accuracy in neutral pronoun predictions, while terms like “nurse” had a much lower replacement rate of 30%. Although we did not use supervised fine-tuning, MOBERT’s exposure to gender-neutralized occupational terms during pre-training likely contributed to its improved performance on those terms in masked language modeling tasks. Since masked language modeling relies on contextual co-occurrence rather than explicit supervision, MOBERT likely developed stronger associations between certain occupations and gender-neutral pronouns due to their repeated exposure in the training data. This suggests that expanding a pre-training dataset to include a broader range of occupations and more balanced representation of male- and female-dominated roles could further improve the model’s performance.
Another potential limitation arises from the process of language alteration itself. Although we carefully designed our pipeline to neutralize pronouns only in contexts where the occupational term was the antecedent, there remains a risk that some instances of gendered language with medically significant context may have been inadvertently modified. While we found no evidence of such errors in our testing, further refinements to the pipeline could incorporate more sophisticated contextual analysis to ensure the protection of patient-specific or trial-related information. Additionally, large-scale gender neutralization poses the challenge of maintaining critical semantic distinctions. While replacing a gendered pronoun with ‘they/them/theirs’ often preserves meaning, certain contexts – such as historical citations or patient narratives – could lead to unintended distortions. Our lexicon filtering approach helps mitigate this by restricting modifications to occupational contexts; however, broader applications must carefully handle edge cases where neutralization may introduce ambiguity or alter medically relevant details. For example, in a PubMed abstract (PMID: 25549443) there is a sentence discussing gender dynamics in surgery: “Suggestions include a change [in] the relationship between a female surgeon and her partner, a supplement of surgeons so that hospitals could change the traditional system of surgery.” Indiscriminate neutralization of “her” to “their” could obscure the focus on challenges specific to female surgeons, weakening the text’s emphasis on gendered professional and personal expectations. Addressing these concerns will require further refinement, including human evaluation studies, dependency parsing for syntactic precision, and additional lexicon constraints to safeguard against unintended language shifts.
@@ Acknowledgements
Elizabeth Schaefer was supported by the Cancer Prevention and Research Institute of Texas (CPRIT RP210045) as part of the Biomedical Informatics, Genomics and Translational Cancer Research Training Program. Kirk Roberts was supported by the National Institutes of Health (R01LM014508, R01LM011934). The authors acknowledge Dhruv Kumar of the University of Texas for serving as a second annotator in the data labeling process, contributing to the double-annotation and reconciliation of the corpus.